OpenStudy (ksaimouli):

determine the equation of a line (2,3) and the midpoint of the segment from (-1,4) to (3,2)

OpenStudy (anonymous):

Get first the midpoint of the line segment. x = (-1 + 3)/2 = 1 y = (4 + 2)/2 = 3 so the midpoint is (1,3). Now get the slope of the line we are looking for using points (2,3) and (1,3). m = (y2 -y1)/(x2 -x1) = (3-3)/(2 -1) = 0 This means that the line is horizontal. By using y = mx+b and knowing that m = 0, the equation of the line is y = 3.